OpenStudy (anonymous):

how do u make 4(2/3) / 7 into a fraction

OpenStudy (australopithecus):

it is a fraction assuming this is the fraction you have \[\frac{4\frac{2}{3}}{7}\]

OpenStudy (australopithecus):

if you want to simplify it I can show you how

OpenStudy (anonymous):

i have to divide 4(2/3) and 7 to get a fraction in simplest form

OpenStudy (anonymous):

i got a decimal that said 2.66666666667

OpenStudy (australopithecus):

\[\frac{4\frac{2}{3}}{7} = \frac{\frac{(4)2}{3}}{7} = \frac{\frac{8}{3}}{7} = \frac{\frac{8}{3}}{\frac{7}{1}}=\frac{8}{3}\frac{1}{7}= \frac{8(1)}{3(7)} = \frac{8}{21}\]

OpenStudy (anonymous):

so the answer is 8/21

OpenStudy (australopithecus):

https://www.wolframalpha.com/input/?i=%284%282%2F3%29%29%2F7

OpenStudy (australopithecus):

do you follow my math?

OpenStudy (australopithecus):

remember any number can be expressed as a fraction using 1 as the demoniator

OpenStudy (anonymous):

yea

OpenStudy (australopithecus):

and fraction multiplication is just as simple as multiplying the numerator by the numerator an denominator by the denominator, anything multiplied by 1 is just the same

OpenStudy (australopithecus):

for division of fractions you just flip the fraction

OpenStudy (anonymous):

ow now i know what u mean and thanks for ur help

OpenStudy (australopithecus):

you flip the bottom fraction and multiply blah

OpenStudy (australopithecus):

well I'm glad you follow, if you are confused about anything just ask I dont mind explaining

OpenStudy (anonymous):

k thanks

OpenStudy (australopithecus):

this is an intermediary step in the problem that I didnt show but might clear up what I'm saying \[\frac{(\frac{4}{1})\frac{2}{3}}{7}\]